OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (7, 4); Slope: 6 A. y – 7 = 6(x + 4) B. y – 4 = 6(x – 7) C. y – 4 = 6(x + 7) D. y + 4 = 6(x – 7)

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) m=slope (x1,y1)=point for this problem m=6 x1=7 y1=4 just plug those three things into the equation and you've got it